Grade 12th passMechanics

a ball is dropped freely while another is thrown vertically downward with an initial velocity v from the same point simultaneously.after t second they are separated by a distance of....solution

Srihari
8 Years agoGrade 12th pass

2 Answers

Soumya Ranjan Mohanty
8 Years ago
For the dropped ball, initial velocity is 0,time=t s and acceleration=g ms^-2 downwards. Distance travelled for dropped ball=x1=ut+1/2 gt^2=0+1/2gt^2.
Similarly, for the ball thrown, initial velocity=v ms^-1, time= t s and acceleration=g ms^-2 downwards. Distance travelled for ball thrown with velocity v=x2=ut+1/2 gt^2=vt+1/2 gt^2.
So distance of separation between them=x2-x1=vt+1/2 gt^2-1/2 gt^2=vt,
therefore, distance of separation between the 2 balls is vt unit distance.
 
Thanks,
Soumya Ranjan Mohanty.
Vikas TU
8 Years ago
Dear Student,

ball1
u=0
v=u+at
=>v=gt

S1=ut+at2/2
=>S1=gt2/2

ball2
u=v
vf=u+at
=>vf=v+gt

S2=vt+gt2/2

distance of separation = S2-S1
=vt.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)